PS1 IPP Czar Logs for the week 2011.05.30 - 2011.06.05

(Up to PS1 IPP Czar Logs)

Monday : 2011.05.30

  • heather was czar
  • restarted publishing, hoping this would allow mops items to be published. This did nothing
  • re-added the publishing survey tasks in stdscience - this fixed mops. No idea why these were missing - the other tasks loaded fine
  • added M31.jtrp to stdscience, since no one was here to stop me...

Tuesday : 2011.05.31

  • 09:40 Bill just queued the ps_ud labels for cleanup
  • 14:31 Bill investigated failed destreak run 534810. Turned out --diff_id 134846 --skycell_id skycell.2.28 had a corrupt file. Fixed with rundiffskyfile.

Wednesday : 2011.06.01

Bill is czar today

  • 10:00 czartool wasn't updating. Followed the instructions on the Processing wiki page to restart it.

Thursday : 2011.06.02

Bill is czar today

  • 09:10 1 warp skyfile was stuck with fault 3 due to database error. Reverted it by hand. 1 diff was stuck due to corrupt warp file --warp_id 204711 --skycell_id skycell.043. Fixed with runwarpskyfile
  • 09:13 all nightlyscience data is through camera stage. Turned chip processing off to let the nightly warps finish.
  • 09:34 set LAP stacks 274446 and 275036 to state drop. These were unfinished because some of the inputs have been cleaned up. These are 2 cases where the same exposures were included multiple times so dropping them is appropriate.
  • 10:00 distribution's pcontrol was pegged at 100% cpu usage. Restarted pantasks.
  • 10:42 turned chip processing back on (7 3pi warps left)

Friday : 2011.06.03

  • 09:40 STS.2010.b i band is complete. Set the runs to be cleaned up.
  • later started running STS.2009.b
  • 16:55 bill removed M31.jtrp since all pending chip inputs fail because the inputs haven't been burntooled. Turned chip.revert back on

Saturday : 2011.06.04

Sunday : 2011.06.05